# Interactive Expressive Motion Generation Using Dynamic Movement Primitives

### v1

@@ Abstract
Our goal is to enable social robots to interact autonomously with humans in a realistic, engaging, and expressive manner. The 12 Principles of Animation [1] are a well-established framework animators use to create movements that make characters appear convincing, dynamic, and emotionally expressive. This paper proposes a novel approach that leverages Dynamic Movement Primitives (DMPs) to implement key animation principles, providing a learnable, explainable, modulable, online adaptable and composable model for automatic expressive motion generation. DMPs, originally developed for general imitation learning in robotics and grounded in a spring-damper system design, offer mathematical properties that make them particularly suitable for this task. Specifically, they enable modulation of the intensities of individual principles and facilitate the decomposition of complex, expressive motion sequences into learnable and parametrizable primitives. We present the mathematical formulation of the parameterized animation principles and demonstrate the effectiveness of our framework through experiments and application on three robotic platforms with different kinematic configurations, in simulation, on actual robots and in a user study. Our results show that the approach allows for creating diverse and nuanced expressions using a single base model.
@@ I Introduction
    We introduce Dynamic Movement Primitives (DMPs) as a new approach for expressive motion generation, and show that representing motor primitives as stable dynamical systems in DMPs has favorable properties for this task.
@@
    We show how DMPs can be extended to modulate a relevant selection of the Principles of Animation. The experiments in simulation, on three different robots, and a user study demonstrate that the approach allows for effectively creating diverse expressive motion through flexible modulation using a single base model.
@@ II-A Expressive Motion Representation and Generation
For the generation of expressive robot motion, methods range from low-level control to learning techniques using generative AI. See also the survey in [12]. The Principles of Animation have been used in several works: In [13], the researchers selected two animation principles for a robot to show forethought before actions and express a reaction to a task’s outcome. In [14], the authors designed animations based on principles for the iCat robot. In subsequent work, they used these animations to compose sequences from the individual animations [15]. The results in [16] show the importance of the timing animation principle in the robot intent expression. The authors of [17] provide a general overview of how animation principles relate to robotics. They apply a selection of principles to emphasize emotion-related gestures for the EMYS robot and verify the expressivity of the outcome with a user study. They extend their work with exemplary robot animations in [18].
@@
Multiple previous works use learning from demonstration (LfD) to align robotic actions with expressive actions from human experts. Some researchers learn cost functions [23] or high-level plans [24]. Relevant to our work are policy learning approaches that map information spaces like perceptual data, system states, contextual cues, or domain-specific heuristics to action spaces. Policies are advantageous for their capabilities to generate immediate low-level trajectories [25]. In [26], models are trained to map human motion features onto spatial velocities for the robot, which can then be controlled with a proportional controller. Using reinforcement learning, whole-body control policies can be made expressive by setting an expressive goal during training [27]. Generally, as stated by [28], enhancements to LfD can be made by including human social cues.
@@ II-B Dynamic Movement Primitives
Various extensions of the original approach have been developed (see the survey in [36]) and include their ability to represent Cartesian orientations in higher dimensions with coupled degrees of freedom [37], their ability to incorporate via-points [38] or task parameters [39], as well as their online adaptation behavior for tasks such as obstacle avoidance by introducing an additional coupling term [40, 41]. Several works address the issue of smoothly joining a sequence of DMPs [42, 43, 44]. Further extensions are probabilistic movement primitives (ProMPs) using Gaussian-based hierarchical Bayesian models [45] and probabilistic dynamic movement primitives (ProDMPs) [46], which combine ProMPs and DMPs to produce smooth trajectory distributions that adhere to initial conditions and capture higher-order statistics.
DMPs have been used in many applications, most prominently for motor skill learning in humanoids [39, 47, 44] and robot manipulation [37], e.g. for assembly [48], but also for movement representation of self-driving cars [23], aerial [49] or underwater field robots [50].
Applications of DMPs in human-robot interaction focus mainly on physical HRI scenarios. Iori et al. [51] introduce a coupling term to time goal reaching in handover scenarios. The coupling term uses the perceived distance from the handover point to influence the robot’s movements to that point. By joining a language-conditioned semantic model with a control model based on DMPs, the authors in [52] allow for abstractions in a manipulation task with language. In [53], a diffusion model determines the parameters of a ProDMP that the authors use for trajectory generation that meets requirements from robot-assisted surgery regarding versatility and quality.
In contrast to these works that typically rely on hand-crafted animations and modulations, our approach enables an animator to use imitation learning to acquire initial sets of re-usable expressive motion primitives systematically and modulate them flexibly along the well-known principles of animation. Unlike recent deep learning methods, our approach is data-efficient, explainable, and, with its theoretical foundation, naturally suited for complex and online adaptive motor behavior. Regarding the DMP literature, our contribution is, to the best of our knowledge, the first application for social HRI and, in particular, for the task of expressive motion generation.
@@ III-A2 Anticipation $\mathcal{P}_{Ant}$
As anticipation should prepare for important parts of the motion, the application should only consider important dimensions or degrees of freedom. Therefore, inverting accelerations only apply to the $N_{\text{Ant}}$ most important dimensions. We determine importance as the differences between minimum and maximum $y$-values, where the dimension with the highest difference is chosen as the most important. Alternatively, the importance inference can be solved differently based on the user’s given requirements or manually set.
@@ III-A6 Secondary Action $\mathcal{P}_{\text{Sec}}$
The primary action is linked to the source, and the target is the modulated dimension. Inverse relations can be specified with $\delta$, either $1$ or $-1$. For values $p_{\text{Sec}}>0$, the secondary action is executed, with higher values leading to stronger secondary actions. Secondary actions are based on specific design choices and require knowledge of the robot’s kinematic structure. For modulation, we assume suitable joint relationships given that specify source, target, and inverse relations.
@@ III-A7 Follow Through $\mathcal{P}_{\text{Follow}}$
Here target refers to the connected body’s dimension and source is the dimension corresponding to the main body. Inverse relationships can be addressed with $\delta$, either $1$ or $-1$. With values $p_{\text{Follow}}>0$, the follow-through action is executed, where higher values lead to more pronounced movements. Note that the selection of target and source is dependent on the system’s kinematic structure and possible intermediate degrees of freedom (DoFs). Therefore, to apply modulation, the target and source must be in the same kinematic chain, with the source higher in the hierarchy. The orientation of their rotation axes should be nearly identical within a defined threshold. Without a threshold, only perfectly aligned joints are modulated. Intermediate DoFs may need to be adjusted to meet this orientation threshold with specified limits.
@@ IV-C User Study
[…]
@@
[…]
@@
$I_{M}$ $I_{P}$uparrow $\downarrow$ $\uparrow$ $\downarrow$ $\uparrow$ $\downarrow$ $\uparrow$ $\downarrow$ $\uparrow$ $\downarrow$ $\uparrow$ $\downarrow$
high $\scriptstyle\mathbf{0.94}$ 0.060.06$\scriptstyle$ 0.060.06 $\scriptstyle\mathbf{0.79}$ 0.210.21$\scriptstyle$ 0.210.21 $\scriptstyle\mathbf{1.00}$ 0.000.00$\scriptstyle$ 0.000.00 $\scriptstyle\mathbf{1.00}$ 0.000.00$\scriptstyle$ 0.000.00 $\scriptstyle\mathbf{0.94}$ 0.060.06$\scriptstyle$ 0.060.06 $\scriptstyle\mathbf{0.56}$ 0.440.44$\scriptstyle$ 0.440.44
low 1.001.00$\scriptstyle$ 1.001.00 $\scriptstyle\mathbf{1.00}$ 0.260.26$\scriptstyle$ 0.260.26 $\scriptstyle\mathbf{0.74}$ 0.000.00$\scriptstyle$ 0.000.00 $\scriptstyle\mathbf{1.00}$ 0.000.00$\scriptstyle$ 0.000.00 $\scriptstyle\mathbf{1.00}$ 0.030.03$\scriptstyle$ 0.030.03 $\scriptstyle\mathbf{0.97}$ 0.290.29$\scriptstyle$ 0.290.29 $\scriptstyle\mathbf{0.71}$

### v2

@@ Abstract
Our goal is to enable social robots to interact autonomously with humans in a realistic, engaging, and expressive manner. The 12 Principles of Animation are a well-established framework animators use to create movements that make characters appear convincing, dynamic, and emotionally expressive. This paper proposes a novel approach that leverages Dynamic Movement Primitives (DMPs) to implement key animation principles, providing a learnable, explainable, modulable, online adaptable and composable model for automatic expressive motion generation. DMPs, originally developed for general imitation learning in robotics and grounded in a spring-damper system design, offer mathematical properties that make them particularly suitable for this task. Specifically, they enable modulation of the intensities of individual principles and facilitate the decomposition of complex, expressive motion sequences into learnable and parametrizable primitives. We present the mathematical formulation of the parameterized animation principles and demonstrate the effectiveness of our framework through experiments and application on three robotic platforms with different kinematic configurations, in simulation, on actual robots and in a user study. Our results show that the approach allows for creating diverse and nuanced expressions using a single base model.
This paper has been accepted for publication at the 2025 IEEE/RSJ International Conference on Intelligent Robots and Systems (IROS).
@@ I Introduction
    We introduce Dynamic Movement Primitives (DMPs) as a new unifying implementation of the Principles of Animation, and show that the model of motor primitives as stable dynamical systems has favorable properties for the task of expressive motion generation.
@@
    Using DMPs, we encode a relevant selection of principles that can be individually modulated to shape a learned reference motion. The experiments in simulation, on three different robots, and a user study demonstrate that the approach allows for effectively creating diverse expressive motion through intuitive modulation using a single base model.
@@ II-A Expressive Motion Representation and Generation
For the generation of expressive robot motion, methods range from low-level control to learning techniques using generative AI. See also the survey in [12]. The Principles of Animation have been used in several works: In [13], the researchers selected two animation principles for a robot to show forethought before actions and express a reaction to a task’s outcome. In [14], the authors designed animations based on principles for the iCat robot. In subsequent work, they used these animations to compose sequences from the individual animations [15]. The results in [16] show the importance of the timing animation principle in the robot intent expression. The authors of [17] provide a general overview of how animation principles relate to robotics. They apply a selection of principles to emphasize emotion-related gestures for the EMYS robot and verify the expressivity of the outcome with a user study. They extend their work with exemplary robot animations in [18]. In [19], the authors propose a method for intent communication of a UAV by modulating flight trajectories based on three principles, each implemented with an individual mathematical model to adjust relevant motion attributes. They show a positive impact on users with respect to usability, safety, and naturalness.
@@
Multiple previous works use learning from demonstration (LfD) to align robotic actions with expressive actions from human experts. Some researchers learn cost functions [24] or high-level plans [25]. Relevant to our work are policy learning approaches that map information spaces like perceptual data, system states, contextual cues, or domain-specific heuristics to action spaces. In [26], models are trained to map human motion features onto spatial velocities for the robot, which can then be controlled with a proportional controller. Using reinforcement learning, whole-body control policies can be made expressive by setting an expressive goal during training [27]. Generally, as stated by [28], enhancements to LfD can be made by including human social cues.
@@ II-B Dynamic Movement Primitives
Various extensions of the original approach have been developed (see the survey in [36]) and include their ability to represent Cartesian orientations in higher dimensions with coupled degrees of freedom [37], their ability to incorporate via-points [38] or task parameters [39], as well as their online adaptation behavior for tasks such as obstacle avoidance by introducing an additional coupling term [40, 41]. Several works address the issue of smoothly joining a sequence of DMPs [42, 43, 44]. Further extensions are probabilistic movement primitives (ProMPs) using Gaussian-based hierarchical Bayesian models [45] and probabilistic dynamic movement primitives (ProDMPs) [46], which combine ProMPs and DMPs to produce bound and smooth trajectory distributions capturing higher-order statistics.
DMPs have been used in many applications, most prominently for motor skill learning in humanoids [39, 47, 44] and robot manipulation [37], e.g. for assembly [48], but also for movement representation of self-driving cars [24], aerial [49] or underwater field robots [50]. Applications of DMPs in human-robot interaction focus mainly on physical HRI scenarios. Iori et al. [51] introduce a coupling term to time goal reaching in handover scenarios. The coupling term uses the perceived distance from the handover point to influence the robot’s movements to that point. By joining a language-conditioned semantic model with a control model based on DMPs, the authors in [52] allow for abstractions in a manipulation task with language. In [53], a diffusion model determines the parameters of a ProDMP that the authors use for trajectory generation that meets requirements from robot-assisted surgery regarding versatility and quality.
In contrast to the works discussed above, our approach differs as follows: Our implementation of the animation principles is, to the best of our knowledge, the first one based on a unified, physically plausible mathematical model that automatically modulates individual principles. This approach overcomes the limitations of handcrafted modulations [13, 14, 15, 16, 17, 18] and ad-hoc models tailored to specific principles [19]. Moreover, and in the spirit of LfD, our approach leverages imitation learning as a way to acquire reference motion. Unlike recent deep learning methods, our approach is data-efficient, explainable, and naturally suited for complex and online adaptive motor behavior. With respect to the DMP literature, our work is the first to adopt DMPs for social HRI and to the task of expressive motion generation.
@@ III-A2 Anticipation $\mathcal{P}_{Ant}$
As anticipation should prepare for important parts of the motion, the application should only consider important dimensions or degrees of freedom. Therefore, inverting accelerations only apply to the $N_{\text{Ant}}$ most important dimensions. We determine importance as the differences between minimum and maximum $y$-values, where the dimension with the highest difference is chosen as the most important. Alternatively, the importance inference can be solved differently based on the user’s given requirements or can be set manually.
@@ III-A6 Secondary Action $\mathcal{P}_{\text{Sec}}$
The primary action is linked to the source, and the target is the modulated dimension. Inverse relations can be specified with $\delta$, either $1$ or 1-1. For values $p_{\text{Sec}}>0$, the secondary action is executed, with higher values leading to stronger secondary actions. Secondary actions are based on specific design choices and require knowledge of the robot’s kinematic structure. For modulation, we assume suitable joint relationships given that specify source, target, and inverse relations.
@@ III-A7 Follow Through $\mathcal{P}_{\text{Follow}}$
Here target refers to the connected body’s dimension and source is the dimension corresponding to the main body. Inverse relationships can be addressed with $\delta$, either $1$ or 1-1. With values $p_{\text{Follow}}>0$, the follow-through action is executed, where higher values lead to more pronounced movements. Note that the selection of target and source is dependent on the system’s kinematic structure and possible intermediate degrees of freedom (DoFs). Therefore, to apply modulation, the target and source must be in the same kinematic chain, with the source higher in the hierarchy. The orientation of their rotation axes should be nearly identical within a defined threshold. Without a threshold, only perfectly aligned joints are modulated. Intermediate DoFs may need to be adjusted to meet this orientation threshold with specified limits.
@@ IV-C User Study
We found that in all instances, participants could recognize the intended emotional expression above chance with statistical significance ($p<.05$). The data also shows that motion with a certain intended emotional expression can be perceived as multiple emotional expressions. This aligns with the findings in [7] and [60], which suggest that certain emotional expressions share basic motion attributes.
@@
TABLE II shows normalized confusion matrices for individual principles comparing modulation intensities $I_{M}$ versus perceived intensities $I_{P}$. The data was analyzed using one-sided independent t-tests. As can be seen from the table, the collected data supports both hypotheses with statistical significance ($p<.05$). This confirms that the scales of the parameters for individual principles follow the intended intuition, which is particularly important as the parameters serve as the main tuning factor for using our method.
@@
$I_{M}$ $I_{P}$ $\uparrow$ $\downarrow$ $\uparrow$ $\downarrow$ $\uparrow$ $\downarrow$ $\uparrow$ $\downarrow$ $\uparrow$ $\downarrow$ $\uparrow$ $\downarrow$
high $\scriptstyle\mathbf{0.94}$ $\scriptstyle 0.06$ $\scriptstyle\mathbf{0.79}$ $\scriptstyle 0.21$ $\scriptstyle\mathbf{1.00}$ $\scriptstyle 0.00$ $\scriptstyle\mathbf{1.00}$ $\scriptstyle 0.00$ $\scriptstyle\mathbf{0.94}$ $\scriptstyle 0.06$ $\scriptstyle\mathbf{0.56}$ $\scriptstyle 0.44$
low $\scriptstyle 1.00$ $\scriptstyle\mathbf{1.00}$ $\scriptstyle 0.26$ $\scriptstyle\mathbf{0.74}$ $\scriptstyle 0.00$ $\scriptstyle\mathbf{1.00}$ $\scriptstyle 0.00$ $\scriptstyle\mathbf{1.00}$ $\scriptstyle 0.03$ $\scriptstyle\mathbf{0.97}$ $\scriptstyle 0.29$ $\scriptstyle\mathbf{0.71}$